Consider the following Gibbs energies at 25 degree
Answer #1

a)

dGrxn for AGcl

AgCl(s) <-> Ag+(Aq) + Cl-(aq)

Grxn = 77.1 -131.2 - (-109.8) =55.7 kJ/mol

dG = -RT*ln(K)

K = exp(-dG/(RT))

K = exp(-55700/(8.31*298)) =1.704*10^-10

Ksp = 1.7*10^-10

c)

For AgBr

AgBr(s) <->Ag+ + Br-

dG = (77.1 -104) - (-96.9) = 70 kJ/mol

d)

for K

dG = -RT*ln(K)

K = exp(-dG/(RT))

K =exp(-70000/(8.314*298) =5.366*10^-13